OpenStudy (anonymous):

9 3/4in multiply 4 3/4 multiply 1/4

OpenStudy (anonymous):

9 3/4 = 39/4 4 3/4 = 19/4 (39/4)(19/4)(1/4) = 741/64 = 11.578125

OpenStudy (mathmale):

I agree with Mel that those mixed numbers (such as 9 3/4 and 4 3/4) should be converted to improper fractions (such as 39/4 and 19/4).\[\frac{ 39 }{ 4 }*\frac{ 19 }{ 4 }*\frac{ 1 }{ 4 }\]